OpenStudy (anonymous):

an and medal Solve for x: -3|x + 7| = -12 x = 5 over 3, x = −19 over 3 x = −3, x = −11 x = −3, x = 11 No solution

OpenStudy (texaschic101):

-3 | x + 7 | = -12 -- divide by -3 | x + 7 | = -12/-3 | x + 7 | = 4 x + 7 = 4 -(x + 7) = 4 x = 4 - 7 -x - 7 = 4 x = -3 -x = 7 + 4 -x = 11 x = -11

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

ily ty.

OpenStudy (texaschic101):

YW :)